Evaluate and classify the critical point of the function f(x, y) = xy


1
Expert's answer
2021-08-09T14:50:51-0400

Given f(x,y) = xy

then fx = y and fy = x

for critical points, fx = 0 and fy = 0

"\implies" x = 0 and y = 0

therefore the only critical point is (0,0)

we know that fx = y and fy = x

"\implies" fxx = 0, fyy = 0 and fxy = 1

Hence, fxx"\cdot" fyy - ( fxy )2 = 0 - 1 = -1 < 0

(0,0) is a saddle point.